2m + n = 0
m + 2n = 3

Found 2 solutions by mahikab, Edwin McCravy:
Answer by mahikab(11)

There are several ways you can solve this problem if you're trying to solve for m and n. You can substitute, or systems of equations. However, I'm going to use substitution:
2m + n = 0 => n = -2m
We can input that in for the other equation:
m + 2n = 3 now becomes: m + 2(-2m) = 3
Now we can solve:
m + 2(−2m) = 3
m + −4m = 3
(m + −4m) = 3 (Combine Like Terms)
−3m=3
m = -1
Now we can input that value in to solve for n:
We said that n = -2m, and m = -1, so n = -2(-1):
n = 2
Your final answer is m = -1, and n = 2, which can also be written as (m,n) = (-1,2)
----------------------
However, if you were solving for m+n:
You would add the two equations!:
2m + n = 0
m + 2n = 3
----------
3m + 3n = 3
Now, you can take 3 common:
3(m+n) = 3
m + n = 1
Your final answer for what m + n equals 1!

Answer by Edwin McCravy(20055)

Oh crap. The other guy did it for you!!!!
Instead of doing your problem for you,  I'll do one exactly like it that you
can use as a step-by-step model to solve yours by.  The one I'll do is:

3m + n = 0
m + 3n = 16
Solve 3m + n = 0 for n
           n = -3m

Substitute (-3m) for n in

    m + 3n = 16
m + 3(-3m) = 16
    m - 9m = 16
       -8m = 16
         m = -2

Substitute -2 for m in
   
         n = -3m
         n = -3(-2)
         n = 6

Now do your problem exactly the same way.

Edwin